OpenStudy (anonymous):

THE RATE OF REACTION IS DOUBLED FOR EVERY 1Oo rise in temprature. the increase in reaction rate as a result of temprature rise from 10o to 100o is 112 512 400 614

OpenStudy (anonymous):

First we need to determine how many times the rate of reaction increases: \[\frac{1000-100}{100}=9\] It doubles every time so we take 2 and raise it to that power: \[2^9=512\]